OpenStudy (anonymous):

Linear Algebra: Let T : P2 => P2 be the linear transform defined by: T(a+b*x+c*x^2) = b+2c*x Find the matrix A of T with respect to the standard basis B = [1, x, x2]. Hint: You might use the fact that P2 is isomorphic to R3.

OpenStudy (anonymous):

so it differentiates, thats fairly easy

OpenStudy (anonymous):

I dont know how to do it though lol

OpenStudy (rosey):

i hate this stuff. u should b a fan of me if u can :)

OpenStudy (anonymous):

you look at the way the function acts on the basic vectors

OpenStudy (anonymous):

T(1) = 0 , which corresponds to a cordinate vector of (0,0,0)

OpenStudy (anonymous):

T(1) = (0, 0, 0) T(x) = (1, 0, 0) T(x^2) = (0, 2, 0)

OpenStudy (anonymous):

^yeh that

OpenStudy (anonymous):

was that row reduced?

OpenStudy (anonymous):

Then setup a matrix using the columns written in the standard basis.

OpenStudy (anonymous):

then you have the first column of the matrix being the first vector

OpenStudy (anonymous):

No, no row reduction in this problem.

OpenStudy (anonymous):

and so on

OpenStudy (anonymous):

\[\left[\begin{matrix}0 & 1 & 0 \\ 0 & 0 & 2 \\ 0 & 0 & 0\end{matrix}\right]\]

OpenStudy (anonymous):

and that final result would be the matrix A of T?

OpenStudy (anonymous):

The matrix A for the transformation T written with respect to the standard basis.

OpenStudy (anonymous):

ahh ok very neat, just curious about one last thing. How did you form this: T(1) = (0, 0, 0) T(x) = (1, 0, 0) T(x^2) = (0, 2, 0) i know you're plugging in the bases of B into T but im not entirely sure whats happening inbetween

OpenStudy (anonymous):

Plug each vector into the transformation and then write the resulting vector out with respect to the standard basis.

OpenStudy (anonymous):

*Plug each basis vector

OpenStudy (anonymous):

T(1) = 0*(1) + 0*(x) + 0*(x^2) T(x) = 1*(1) + 0*(x) + 0*(x^2) T(x^2) = 0*(1) + 2*(x) + 0*(x^2)

OpenStudy (anonymous):

what i dont understand is how you got 000 100 020

OpenStudy (anonymous):

Plug in the first basis element into the transformation. What do you get?

OpenStudy (anonymous):

T(a+b*1+c*1^2) = b+2c*1?

OpenStudy (anonymous):

Plug in the first basis element. The first basis element is 1.

OpenStudy (anonymous):

i know i replaced x with 1

OpenStudy (anonymous):

No, that's wrong. x is a place holder, these are vectors. You could just as well be dealing with a basis of apples and oranges.

OpenStudy (anonymous):

This is the mapping [a, b, c] = [b, 2c, 0]

OpenStudy (anonymous):

Mapping: [a, b, c] = [b, 2c, 0] [1, 0, 0] = [0, 0, 0] [0, 1, 0] = [1, 0, 0] [0, 0, 1] = [0, 2, 0]

OpenStudy (anonymous):

oh i see shouldnt it be [a, b, c] = [0, b, 2c]? though seeing a would go first or if a vector is empty does it go at the end of the mapped matrix

OpenStudy (anonymous):

no, does anything get mapped to x^2?

OpenStudy (anonymous):

ahhh i see now

OpenStudy (anonymous):

ok i see how you arrived at the mapping for the beginning, but not this part [1, 0, 0] = [0, 0, 0] [0, 1, 0] = [1, 0, 0] [0, 0, 1] = [0, 2, 0] did you take the basis and put it with [a, b, c] = [b, 2c, 0] to arrive at the final result above?

OpenStudy (anonymous):

A matrix is really just an ordered set of basis vectors that have been transformed and written with respect to the basis for the target space.

OpenStudy (anonymous):

So perform the mapping for each basis vector and write it with respect to the standard basis.

OpenStudy (anonymous):

i still dont get how you got [1, 0, 0] = [0, 0, 0] [0, 1, 0] = [1, 0, 0] [0, 0, 1] = [0, 2, 0]

OpenStudy (anonymous):

There is a bijection between any vector space and F^n for any field F. In this case we consider the field R. The point is you can perform a bijection between any vector in a given vector space and the coefficients in the linear combination.

OpenStudy (anonymous):

1, becomes (1, 0, 0) x, becomes (0, 1, 0) x^2, becomes (0, 0, 1)

OpenStudy (anonymous):

plug each standard basis element into the map [a, b, c] = [b, 2c, 0] [1, 0, 0] = [0, 0, 0] [0, 1, 0] = [1, 0, 0] [0, 0, 1] = [0, 2, 0]

OpenStudy (anonymous):

ahh i see i see! thank you very much Alchemista again for your help today XD